### v1

@@ Abstract
The proof of this theorem can be essentially deduced from the results of [AuHa14b, AretE15], see §6.1.
@@
  • Approach I: Finite energy spaces in dimensions $n\geq 3$. Solve the exterior Steklov problem using variational characterisation ($\mathrm{FE}$) in the space of functions of finite energy, i.e., functions with their gradients in $L^{2}\left(\Omega^{\mathrm{ext}}\right)$, and satisfying a certain decay condition at infinity, see §3.1.2. This approach was developed in [AuHa14b, AretE15].
@@
  • Approach III: Truncated domains. Given $R>R_{0}(\Omega)$, consider a mixed Steklov–Dirichlet eigenvalue problem ($\mathrm{SD}_{R}$) on $\Omega^{\mathrm{ext}}_{R}$ with Steklov condition on $\partial\Omega$ and Dirichlet condition on $\partial B_{R}$, and take the limit of eigenvalues and eigenfunctions as $R\to\infty$, see §3.2. This approach was explored in [AretE15] in dimension $n\geq 3$.
@@
The equivalence of these formulations is understood as follows. We will show that for the Approaches I, II, and V, the corresponding eigenvalues and eigenfunctions coincide with those of the exterior Steklov problem ($\mathrm{ES}$), see Definition 1.2. For approaches III and IV, we will prove that the $k$th eigenvalues of the problems ($\mathrm{SD}_{R}$) as $R\to\infty$ and ($\mathrm{H}_{\Lambda}$) as $\Lambda\searrow 0$ converge to $\sigma_{k}\left(\Omega^{\mathrm{ext}}\right)$ for all $k\geq 1$, and the corresponding eigenfunctions converge in an appropriate sense to the eigenfunctions $u_{k}$ of the exterior Steklov problem. For schematic relation between different approaches, see Figures 2 and 3. For the precise statements, we refer to §5.
@@
There exist other possible approaches to the exterior Steklov problem. In particular, one can modify the truncated domains approach and impose Neumann conditions on $\partial B_{R}$ instead of the Dirichlet conditions. In two dimensions this formulation is equivalent to the others, see §5.2. At the same time, as was shown in [AretE15], in dimensions $n\geq 3$ the Neumann truncation does not provide an equivalent formulation to the ones described above. In fact, this can be seen from the fact that in dimensions $n\geq 3$ the first exterior Steklov eigenvalue is positive, while the Neumann truncation yields a zero eigenvalue. In these dimensions, the Neumann truncation corresponds instead to the exterior Steklov problem with a decay condition on the gradient (vanishing flow), as described in §4.3.2.
A problem related to ($\mathrm{H}_{\Lambda}$) has been studied in [HelNi26]. Instead of introducing a Helmholtz parameter $\Lambda$, one can consider a magnetic Laplacian and examine the spectral behaviour as the magnetic potential vanishes. It would be interesting to verify whether in the limit one gets the same exterior Steklov problem as above. ∎
FiniteenergyTruncatedDirichletTruncatedNeumannHelmholtzequationLayerPotentialsOur ExternalSteklovVanishingflowTheorem 5.4Theorem 5.6Corollary 4.6Proof of Theorem 1.3Theorem 5.8
@@
ConformalmappingTruncatedDirichletTruncatedNeumannHelmholtzequationLayerPotentialsOur ExternalSteklovVanishingflowTheorem 5.9Theorem 5.13Theorem 5.15Proof of Theorem 1.3Theorem 5.10Remark 4.9
@@ 1.3.1 Courant’s theorem and multiplicity bounds
Recall that a nodal domain of a continuous function is a connected component of the complement of its nodal set (that is, the set of zero values of the function). One of the fundamental properties of the Laplace and the (interior) Steklov eigenfunctions is Courant’s domain theorem. It is easy to show that it holds for the exterior Steklov eigenfunctions as well.
@@
Similarly to the interior problem, there is no direct analogue of Courant’s theorem for the nodal domains of the exterior Dirichlet-to-Neumann map (i.e., for the boundary nodal domains of Steklov eigenfunctions). In two dimensions, Theorem 1.8 and elementary topological considerations yield a bound on the number of boundary sign changes, cf. [GirPo17, Section 6]. Also note that using the results on the exterior Dirichlet-to-Neumann semigroup obtained in [AretE15, Section 5] one can deduce that the first exterior Steklov eigenfunction on any Lipschitz domain can be chosen to be strictly positive on the boundary, cf. [AretEGl20]. ∎
@@ 1.3.2 Escobar-type lower bounds for the first eigenvalue
Let $k\in\mathbb{N}$, and $\alpha_{1},\dots,\alpha_{k}$ be distinct positive real numbers. Their logarithmic mean $L(\alpha_{1},\ldots,\alpha_{k})$ is defined as [Pi85]
$L(\alpha_{1},\ldots,\alpha_{k})=\frac{1}{\sum_{j=1}^{k}\frac{\alpha_{j}^{k-2}}{\prod\limits_{i=1,\,i\neq j}^{k}(\alpha_{j}-\alpha_{i})}\log\alpha_{j}}.$
If the $\alpha_{i}$ are not distinct, $L$ is defined as the limit of the right-hand side.
@@ 1.3.5 Eigenvalue asymptotics
where $r\in(\rho,\infty)$, $\theta\in(-\pi,\pi]$. This series converges absolutely in $B_{\rho}^{\mathrm{ext}}$ and uniformly on any compact subset.
@@
We additionally have
Proposition 2.3.
Consider a disk $B_{\rho}\in\mathbb{R}^{2}$, $\rho>0$, and let $f\in H^{1/2}(\partial B_{\rho})$. Then $\nabla\left(\mathcal{H}^{\mathrm{ext}}f\right)\in L^{2}\left(B_{\rho}^{\mathrm{ext}}\right)$.
Proof.
We expand $f$ in the Fourier series in the boundary traces of eigenfunctions (2.1) on $\partial B_{\rho}$, re-writing them in the complex form, to get
$f=\sum_{m\in\mathbb{Z}}\hat{f}_{m}\rho^{-|m|}\mathrm{e}^{\mathrm{i}m\theta},$
where the condition $f\in H^{1/2}(\partial B_{\rho})$ is equivalent to
$\sum_{m\in\mathbb{Z}}\sqrt{1+|m|^{2}}\rho^{-2|m|}\left|\hat{f}_{m}\right|^{2}<\infty.$ (2.5)
The harmonic extension $u:=\mathcal{H}^{\mathrm{ext}}f$ is then
$u(r,\theta)=\sum_{m\in\mathbb{Z}}\hat{f}_{m}r^{-|m|}\mathrm{e}^{\mathrm{i}m\theta},$
and explicit integration gives
$\left\|\nabla u\right\|^{2}_{L^{2}\left(B_{\rho}^{\mathrm{ext}}\right)}=\sum_{m\in\mathbb{Z}}|m|\rho^{-2|m|}\left|\hat{f}_{m}\right|^{2},$
which is finite by (2.5). ∎
A similar result holds in higher dimensions as well, see Proposition 2.5.
@@
In order to fix the terminology, let us recall that the spectrum of the Laplace–Beltrami operator $\Delta_{\mathbb{S}^{n-1}}$ on the unit sphere $\mathbb{S}^{n-1}\subset\mathbb{R}^{n}$, $n\geq 3$, consists of eigenvalues $\alpha_{\ell}:=\ell(\ell+n-2)$, $\ell\in\{0\}\cup\mathbb{N}$, of multiplicity
@@
The elements of the eigenspace of $\Delta_{\mathbb{S}^{n-1}}$ corresponding to $\alpha_{\ell}$ are known as spherical harmonics of degree $\ell$, and one can choose an orthonormal basis $\left\{Y_{\ell,i}\right\}_{i=1}^{d_{n,\ell}}$ in each such eigenspace.
@@
$u_{(\ell,i)}^{\mathrm{D}}(r,\theta)=\left(r^{2-n-\ell}-R^{2-n-2\ell}r^{\ell}\right){Y}_{\ell,i}(\theta),\qquad\ell\in\{0\}\cup\mathbb{N},\quad i=1,\ldots,d_{n,\ell},$
@@
We also have the direct analogue of Proposition 2.3, which is easily proved in a similar manner.
@@
Consider a ball $B_{\rho}\in\mathbb{R}^{n}$, $\rho>0$, $n\geq 3$, and let $f\in H^{1/2}(\partial B_{\rho})$. Then $\nabla\left(\mathcal{H}^{\mathrm{ext}}f\right)\in L^{2}\left(B_{\rho}^{\mathrm{ext}}\right)$.
@@
For both geometries, the exterior eigenvalues and eigenfunctions are computed using the conformal transform method of §3.4 and the finite element calculations in FreeFEM [Hec12] for the corresponding interior domain problem, see scripts cited at the front page. For an alternative FEM approach to the exterior Steklov problem, see also [GrCh25] and [ChaGr24].
@@
We start with the global approach introduced by Arendt and ter Elst [AretE15] for the Dirichlet-to-Neumann operators, as well as by Auchmuty and Han [AuHa14b] for the weak formulation of the Steklov problem. We work directly in function spaces over the full exterior domain. It is in this setting that we will present the spectral theorem for the exterior problem under consideration. It will be considered as a reference problem, that will later be compared to various other methods. The planar case is different and will be discussed separately.
@@ 3.1.1 Dirichlet-to-Neumann operators
For $n\geq 3$, Arendt and ter Elst [AretE15] introduced two exterior Dirichlet-to-Neumann operators $A^{\mathrm{N}}$ and $A^{\mathrm{D}}$ on $L^{2}(\partial\Omega)$ using the following two function spaces,
@@
is complete, the trace operator $\boldsymbol{\tau}:W\left(\Omega^{\mathrm{ext}}\right)\to L^{2}(\partial\Omega)$, $\boldsymbol{\tau}:u\mapsto u|_{\partial\Omega}$, is compact, and the norm defined above is equivalent to $u\mapsto\left(\|\nabla u\|_{L^{2}\left(\Omega^{\mathrm{ext}}\right)}^{2}+\|u\|_{L^{2}(\partial\Omega)}^{2}\right)^{1/2}$, see [AretE15, Proposition 2.5].
@@
is the self-adjoint operator on $L^{2}(\partial\Omega)$ associated with $\left(W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right),Q,\boldsymbol{\tau}\right)$, see [AretE15]. ∎
The following result follows from [AretE15, Theorem 5.9 and the remark before Proposition 5.3].
@@
Let $P_{0}^{\perp}$ denote the orthogonal projection on the orthogonal complement of $\operatorname{Span}\{{h}_{0}\}$ in $H$, and let $\mathcal{R}_{\mathcal{A}}[{h}]$ and $\mathcal{R}_{\mathcal{B}}[{h}]$ be the Rayleigh quotients of $\mathcal{A}$ and $\mathcal{B}$ evaluated at ${h}\in D\setminus\{0\}$. Fix $k\in\mathbb{N}$, and recall the variational principle
@@
Take any subspace $\mathcal{L}\subset D$ with $\dim\mathcal{L}=k+1$. Then $\dim P_{0}^{\perp}\mathcal{L}\geq k$. Choose now any subspace $\mathcal{M}\subset P_{0}^{\perp}\mathcal{L}$, $\dim\mathcal{M}=k$. Then restrictions of the operators $\mathcal{A}$ and $\mathcal{B}$ to $\mathcal{M}$ coincide, and therefore, as $\mathcal{M}\subset\mathcal{L}$,
@@
Let us now take, in the right- and left-hand sides of (3.3), the minimum over all subspaces $\mathcal{L}\subset D$, $\dim\mathcal{L}=k+1$, and all $\mathcal{M}\subset P_{0}^{\perp}\mathcal{L}$, $\dim\mathcal{M}=k$. As the right-hand side of (3.3) is independent of $\mathcal{M}$, we get
$\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\\ \mathcal{M}\subset P_{0}^{\perp}\mathcal{L}\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{L}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{A}}[{h}]=\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{L}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{A}}[{h}]=\lambda_{k+1}(\mathcal{A}).$
@@
$\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\\ \mathcal{M}\subset P_{0}^{\perp}\mathcal{L}\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{M}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{B}}[{h}]\geq\min_{\begin{subarray}{c}\mathcal{M}\subset D\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{M}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{B}}[{h}]=\lambda_{k}(\mathcal{B}),$

### v2

@@ Abstract
The proof of this theorem can be essentially deduced from the results of [AuHa14b, ArtE15], see §6.1.
@@
  • Approach I: Finite energy spaces in dimensions $n\geq 3$. Solve the exterior Steklov problem using variational characterisation ($\mathrm{FE}$) in the space of functions of finite energy, i.e., functions with their gradients in $L^{2}\left(\Omega^{\mathrm{ext}}\right)$, and satisfying a certain decay condition at infinity, see §3.1.2. This approach was developed in [AuHa14b, ArtE15].
@@
  • Approach III: Truncated domains. Given $R>R_{0}(\Omega)$, consider a mixed Steklov–Dirichlet eigenvalue problem ($\mathrm{SD}_{R}$) on $\Omega^{\mathrm{ext}}_{R}$ with Steklov condition on $\partial\Omega$ and Dirichlet condition on $\partial B_{R}$, and take the limit of eigenvalues and eigenfunctions as $R\to\infty$, see §3.2. This approach was explored in [ArtE15] in dimension $n\geq 3$.
@@
The equivalence of these formulations is understood as follows. We will show that for the Approaches I, II, and V, the corresponding eigenvalues and eigenfunctions coincide with those of the exterior Steklov problem ($\mathrm{ES}$), see Definition 1.2. For approaches III and IV, we will prove that the $k$th eigenvalues of the problems ($\mathrm{SD}_{R}$) as $R\to\infty$ and ($\mathrm{H}_{\Lambda}$) as $\Lambda\searrow 0$ converge to $\sigma_{k}\left(\Omega^{\mathrm{ext}}\right)$ for all $k\geq 1$, and the corresponding eigenfunctions converge in an appropriate sense to the eigenfunctions $u_{k}$ of the exterior Steklov problem. For schematic relation between different approaches, see Figures 2 and 3. For the precise statements, we refer to §4 and §5.
@@
There exist other possible approaches to the exterior Steklov problem. One approach, useful for numerical analysis, is to reduce ($\mathrm{ES}$) to a mixed problem ($\mathrm{S}\Psi_{\rho}$) in a bounded domain $\Omega^{\mathrm{ext}}_{\rho}$, with the Steklov condition on $\partial\Omega$ and a pseudodifferential matching condition on $\partial B_{\rho}$, see §2.4.
Also, one can modify the truncated domains approach and impose Neumann conditions on $\partial B_{R}$ instead of the Dirichlet conditions. In two dimensions this formulation is equivalent to the others, see §5.2. At the same time, as was shown in [ArtE15], in dimensions $n\geq 3$ the Neumann truncation does not provide an equivalent formulation to the ones described above. In fact, this can be seen from the fact that in dimensions $n\geq 3$ the first exterior Steklov eigenvalue is positive, while the Neumann truncation yields a zero eigenvalue. In these dimensions, the Neumann truncation corresponds instead to the exterior Steklov problem with a decay condition on the gradient (vanishing flow), as described in §4.3.2.
An approach somewhat related to ($\mathrm{H}_{\Lambda}$) has been introduced in [HelNi26], see also [HelKaNi25]. Instead of introducing a Helmholtz parameter $\Lambda$, one can consider a magnetic Laplacian and examine the spectral behaviour as the magnetic potential vanishes. It would be interesting to verify whether in the limit one gets the same exterior Steklov problem as above. ∎
FiniteenergyTruncatedDirichletTruncatedNeumannHelmholtzequationLayerpotentialsExteriorSteklovVanishingflowTheorem 5.4Theorem 5.6Corollary 4.6Proof of Theorem 1.3Theorem 5.8
@@
ConformalmappingTruncatedDirichletTruncatedNeumannHelmholtzequationLayerpotentialsExteriorSteklovVanishingflowTheorem 5.9Theorem 5.14Theorem 5.16Proof of Theorem 1.3Theorem 5.10Remark 4.9
@@ 1.3.1 Courant’s theorem and multiplicity bounds
Recall that a nodal domain of a continuous function is a connected component of the complement of its nodal set (that is, the set of zero values of the function). One of the fundamental properties of the Laplace and the (interior) Steklov eigenfunctions is Courant’s nodal domain theorem. It is easy to show that it holds for the exterior Steklov eigenfunctions as well.
@@
Similarly to the interior problem, there is no direct analogue of Courant’s theorem for the nodal domains of the eigenfunctions of the exterior Dirichlet-to-Neumann map (i.e., for the boundary nodal domains of Steklov eigenfunctions). In two dimensions, Theorem 1.8 and elementary topological considerations yield a bound on the number of boundary sign changes, cf. [GirPo17, Section 6]. Also note that using the results on the exterior Dirichlet-to-Neumann semigroup obtained in [ArtE15, Section 5] one can deduce that the first exterior Steklov eigenfunction on any Lipschitz domain can be chosen to be strictly positive on the boundary, cf. [ArtEGl20]. ∎
@@ 1.3.2 Escobar-type lower bounds for the first eigenvalue
Let $k\in\mathbb{N}$, and $\alpha_{1},\dots,\alpha_{k}$ be distinct positive real numbers. Following [Pi85], we define their logarithmic mean $L(\alpha_{1},\ldots,\alpha_{k})$ as
$L(\alpha_{1},\ldots,\alpha_{k})=\frac{1}{(k-1)\sum_{j=1}^{k}\frac{\alpha_{j}^{k-2}}{\prod\limits_{i=1,\,i\neq j}^{k}(\alpha_{j}-\alpha_{i})}\log\alpha_{j}}.$ (1.3)
If the $\alpha_{i}$ are not distinct, or if an $\alpha_{i}$ vanishes, $L$ is defined as the limit of the right-hand side of (1.3). Note that for $k>2$ there exist other definitions of the logarithmic mean in the literature. For $k=2$, however, they all agree and yield
$L(\alpha_{1},\alpha_{2})=\begin{cases}\frac{\alpha_{1}-\alpha_{2}}{\log\alpha_{1}-\log\alpha_{2}}&\text{ for }\alpha_{1}\neq\alpha_{2},\\ \alpha_{1}&\text{ for }\alpha_{1}=\alpha_{2}.\end{cases}$
@@ 1.3.5 Eigenvalue asymptotics
where $r\in(\rho,\infty)$, $\theta\in\mathbb{S}^{1}$. This series converges absolutely in $B_{\rho}^{\mathrm{ext}}$ and uniformly on any compact subset.
@@
In order to fix the terminology, let us recall that the spectrum of the Laplace–Beltrami operator $\Delta_{\mathbb{S}^{n-1}}$ on the unit sphere $\mathbb{S}^{n-1}\subset\mathbb{R}^{n}$, $n\geq 3$, consists of eigenvalues $\alpha_{(\ell)}:=\ell(\ell+n-2)$, $\ell\in\{0\}\cup\mathbb{N}$, of multiplicity
@@
The elements of the eigenspace of $\Delta_{\mathbb{S}^{n-1}}$ corresponding to $\alpha_{(\ell)}$ are known as spherical harmonics of degree $\ell$, and one can choose an $L^{2}\left(\mathbb{S}^{n-1}\right)$-orthonormal basis $\left\{Y_{\ell,i}\right\}_{i=1}^{d_{n,\ell}}$ in each such eigenspace.
@@
$u_{(\ell,i)}^{\mathrm{D}}(r,\theta)=\left(r^{2-n-\ell}-R^{2-n-2\ell}r^{\ell}\right){Y}_{\ell,i}(\theta),\qquad\ell\in\{0\}\cup\mathbb{N},\quad i=1,\ldots,d_{n,\ell}.$
@@
§2.3.  Exterior Dirichlet-to-Neumann maps and harmonic extensions for disks and balls
It is easily seen that for any $n\geq 2$, the functions
$\widetilde{Y}_{\ell,i}(\theta):=\rho^{(1-n)/2}Y_{\ell,i}(\theta),\qquad\ell\in\mathbb{N},\quad i=1,\ldots,d_{n,\ell},$
form an $L^{2}\left(\partial B_{\rho}\right)$-orthonormal basis of eigenfunctions of the exterior Dirichlet-to-Neumann map $\mathcal{D}^{\mathrm{ext}}$ on $\partial B_{\rho}$, where for uniformity we set in the planar case
$Y_{0,1}(\theta):=\frac{1}{\sqrt{2\pi}},\qquad Y_{\ell,1}(\theta):=\frac{1}{\sqrt{\pi}}\sin(\ell\theta),\quad Y_{\ell,2}(\theta):=\frac{1}{\sqrt{\pi}}\cos(\ell\theta),\qquad\ell\in\mathbb{N}.$
Therefore, $\mathcal{D}^{\mathrm{ext}}$ acts on any $f\in H^{\frac{1}{2}}\left(\partial B_{\rho}\right)$ as
$\mathcal{D}^{\mathrm{ext}}f=\sum_{\ell=0}^{\infty}\sum_{i=1}^{d_{n,\ell}}\sigma_{(\ell)}\left(B_{\rho}^{\mathrm{ext}}\right)\widehat{f}_{\ell,i}\widetilde{Y}_{\ell,i}\in H^{-\frac{1}{2}}\left(\partial B_{\rho}\right),\qquad\text{where}\quad\widehat{f}_{\ell,i}:=\left(f,\widetilde{Y}_{\ell,i}\right)_{L^{2}\left(\partial B_{\rho}\right)}.$ (2.7)
We additionally have
@@
Consider, for $n\geq 2$, a ball $B_{\rho}\in\mathbb{R}^{n}$, $\rho>0$, and let $f\in H^{1/2}\left(\partial B_{\rho}\right)$. Then $\nabla\left(\mathcal{H}^{\mathrm{ext}}f\right)\in L^{2}\left(B_{\rho}^{\mathrm{ext}}\right)$.
Proof.
We expand $f$ in the Fourier series in the basis of spherical harmonics on $\partial B_{\rho}$, to get
$f=\sum_{\ell=0}^{\infty}\sum_{i=1}^{d_{n,\ell}}\widehat{f}_{\ell,i}\widetilde{Y}_{\ell,i},$
where the condition $f\in H^{1/2}\left(\partial B_{\rho}\right)$ is equivalent to
$\sum_{\ell=0}^{\infty}\sum_{i=1}^{d_{n,\ell}}\sqrt{1+\frac{\alpha_{(\ell)}}{\rho^{2}}}\left|\widehat{f}_{\ell,i}\right|^{2}<\infty.$ (2.8)
The harmonic extension $u:=\mathcal{H}^{\mathrm{ext}}f$ is then
$u(r,\theta)=\sum_{\ell=0}^{\infty}\sum_{i=1}^{d_{n,\ell}}\widehat{f}_{\ell,i}r^{2-n-\ell}\widetilde{Y}_{\ell,i},$
and explicit integration gives
$\left\|\nabla u\right\|^{2}_{L^{2}\left(B_{\rho}^{\mathrm{ext}}\right)}=\sum_{\ell=0}^{\infty}\sum_{i=1}^{d_{n,\ell}}\sigma_{(\ell)}\left(B_{\rho}^{\mathrm{ext}}\right)\left|\widehat{f}_{\ell,i}\right|^{2},$ (2.9)
which is finite by (2.8) due to the fact that
$\lim_{\ell\to\infty}\frac{\sigma_{(\ell)}\left(B_{\rho}^{\mathrm{ext}}\right)}{\sqrt{1+\frac{\alpha_{(\ell)}}{\rho^{2}}}}=1.$
@@
For both geometries, the exterior eigenvalues and eigenfunctions are computed using the conformal transform method of §3.4 and the finite element calculations in FreeFEM [Hec12] for the corresponding interior domain problem, see scripts cited at the front page.
An alternative numerical approach to the exterior Steklov problem, see, e.g., [GrCh25, ChaGr24], can be realised via the domain decomposition technique by reducing the problem to a mixed Steklov–matching condition problem in a bounded domain. Let $\Omega\subset\mathbb{R}^{n}$ be a bounded open set with Lipschitz boundary and connected $\Omega^{\mathrm{ext}}$, and fix any $\rho>R_{0}(\Omega)$. Then it is easy to see that the problem ($\mathrm{ES}$) is equivalent to
$\begin{cases}\Delta u=0\qquad&\text{ in }\Omega^{\mathrm{ext}}_{\rho},\\ \partial_{\nu}u=\sigma u\qquad&\text{ on }\partial\Omega,\\ \left(\partial_{\overline{\nu}}+\mathcal{D}^{\mathrm{ext}}\right)u=0\qquad&\text{ on }\partial B_{\rho},\end{cases}$ ($\mathrm{S}\Psi_{\rho}$)
where the exterior Dirichlet-to-Neumann map $\mathcal{D}^{\mathrm{ext}}$ appearing in the pseudodifferential boundary condition on $\partial B_{\rho}$ is given by (2.7). In addition to a possible numerical implementation, the problem ($\mathrm{S}\Psi_{\rho}$) can be viewed as yet another equivalent formulation of the exterior Steklov problem.
@@
We start with the global approach introduced by Arendt and ter Elst [ArtE15] for the Dirichlet-to-Neumann operators, as well as by Auchmuty and Han [AuHa14b] for the weak formulation of the Steklov problem. We work directly in function spaces over the full exterior domain. It is in this setting that we will present the spectral theorem for the exterior problem under consideration. It will be considered as a reference problem, that will later be compared to various other methods. The planar case is different and will be discussed separately.
@@ 3.1.1 Dirichlet-to-Neumann operators
For $n\geq 3$, Arendt and ter Elst [ArtE15] introduced two exterior Dirichlet-to-Neumann operators $A^{\mathrm{N}}$ and $A^{\mathrm{D}}$ on $L^{2}(\partial\Omega)$ using the following two function spaces,
@@
is complete, the trace operator $\boldsymbol{\tau}:W\left(\Omega^{\mathrm{ext}}\right)\to L^{2}(\partial\Omega)$, $\boldsymbol{\tau}:u\mapsto u|_{\partial\Omega}$, is compact, and the norm defined above is equivalent to $u\mapsto\left(\|\nabla u\|_{L^{2}\left(\Omega^{\mathrm{ext}}\right)}^{2}+\|u\|_{L^{2}(\partial\Omega)}^{2}\right)^{1/2}$, see [ArtE15, Proposition 2.5].
@@
is the self-adjoint operator on $L^{2}(\partial\Omega)$ associated with $\left(W^{\mathrm{D}}\left(\Omega^{\mathrm{ext}}\right),Q,\boldsymbol{\tau}\right)$, see [ArtE15]. ∎
The following result follows from [ArtE15, Theorem 5.9 and the remark before Proposition 5.3].
@@
Let $\mathcal{R}_{\mathcal{A}}[{h}]$ and $\mathcal{R}_{\mathcal{B}}[{h}]$ be the Rayleigh quotients of $\mathcal{A}$ and $\mathcal{B}$ evaluated at ${h}\in D\setminus\{0\}$. Fix $k\in\mathbb{N}$, and recall the variational principle
@@
Take any subspace $\mathcal{L}\subset D$ with $\dim\mathcal{L}=k+1$, and set $\widetilde{\mathcal{L}}:=\left\{h\in\mathcal{L}:\left(h,h_{0}\right)_{H}=0\right\}$. Then $\dim\widetilde{\mathcal{L}}\geq k$. Choose now any subspace $\mathcal{M}\subset\widetilde{\mathcal{L}}$, $\dim\mathcal{M}=k$. Then restrictions of the operators $\mathcal{A}$ and $\mathcal{B}$ to $\mathcal{M}$ coincide, and therefore, as $\mathcal{M}\subset\mathcal{L}$,
@@
Let us now take, in the right- and left-hand sides of (3.3), the minimum over all subspaces $\mathcal{L}\subset D$, $\dim\mathcal{L}=k+1$, and all $\mathcal{M}\subset\widetilde{\mathcal{L}}$, $\dim\mathcal{M}=k$. As the right-hand side of (3.3) is independent of $\mathcal{M}$, we get
$\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\\ \mathcal{M}\subset\widetilde{\mathcal{L}}\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{L}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{A}}[{h}]=\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{L}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{A}}[{h}]=\lambda_{k+1}(\mathcal{A}).$
@@
$\min_{\begin{subarray}{c}\mathcal{L}\subset D\\ \dim\mathcal{L}=k+1\\ \mathcal{M}\subset\widetilde{\mathcal{L}}\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{M}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{B}}[{h}]\geq\min_{\begin{subarray}{c}\mathcal{M}\subset D\\ \dim\mathcal{M}=k\end{subarray}}\ \max_{\begin{subarray}{c}{h}\in\mathcal{M}\\ {h}\neq 0\end{subarray}}\ \mathcal{R}_{\mathcal{B}}[{h}]=\lambda_{k}(\mathcal{B}),$